. (a) How many different positive three-digit whole numbers can be formed from

the four digits 2, 6, 7, and 9 if any digit can be repeated?

(b) How many different positive whole numbers less than 1000 can be formed

from 2, 6, 7, 9 if any digit can be repeated?

(c) How many numbers in part (b) are less than 680 (i.e. up to 679)?

(d) What is the probability that a positive whole number less than 1000, chosen

at random from 2, 6, 7, 9 and allowing any digit to be repeated, will be less

than 680?

7. Answer question 7 again for the case where the digits 2, 6, 7, 9 can not be repeated


1
Expert's answer
2022-03-17T18:57:54-0400

Solution

a) Repetitions are allowed

"n=4, ~r=3"


"=n^r =4^3"

"=64"


(b) Repetitions are allowed

"n=4, ~r=1"

"=n^r=4^1=4"


"n=4, ~r=2"

"=n^r=4^2=16"


"n=4,~r=3"

"=n^r=4^3=64"


Total number that can be formed

"=4+16+64"

"=84"


(c) Repetitions allowed.

1.Numbers from 692 to 699

"692,696,697,699"


2. Numbers from 722 to 799 "n=4, r=2"

"=n^r,=4^2=16"


3. Numbers from 922 to 999

"n=4,~ r=2"

"=n^r=4^2=16"


Less than 680

"84-(4+16+16)=58"


(d) Probability that a number is less than 680

"P(X<680)=\dfrac {58}{84}\approx 0.69"


7. Repetition not allowed

(a) Possible samples

"n=5,r=3"


"=C_4^3= \binom{4}{3}=4" samples


(b) Possible samples less than 1000


since "r=3" all "4" samples will be less than "1000"


(c) Samples less than 680


The samples will be as follows

"(2,6,7)~(2,6,9)~(2,7,9)~(6,7,9)"


All "4" samples will be less than 680.


(d) Probability that a number is less than 680


"P(X<680)=\dfrac44=1"